FIND THE VALUE OF (K) FOR WHICH THERE IS NO SOLUTION OF EQUATIONS
(K+1)X+8Y=4K AND
(K)X+ (K+3)Y=3K-1

Grade:12

1 Answers

bharat bajaj IIT Delhi
10 years ago
If there is no solution,
K+1/K = 8/K+3 = 4K/3K-1
K^2 -4K + 3 = 0
K = 1,3
6K-2 = k^2 + 3K
K^2 - 3K + 2 = 0
K = 2,1
Hence, K = 1 for no solution
Thanks
Bharat Bajaj
IIT Delhi
askiitians faculty